OpenStudy (anonymous):

Integrate 5*csc^4x*cot^6x*dx

OpenStudy (luigi0210):

Hello @meggie94 Welcome to Openstudy :)

OpenStudy (anonymous):

Cheating is not permitted...

OpenStudy (anonymous):

Cheating is against terms and services, if you took the time to read it...

OpenStudy (luigi0210):

@shaagasda Yes, cheating on tests, exams, etc is not permitted

OpenStudy (anonymous):

Tell that to him...

OpenStudy (luigi0210):

What makes you think he's cheating?

OpenStudy (anonymous):

That question was in a math packet I recently did.

OpenStudy (anonymous):

I'm not trying to cheat, I just really don't know what to do...

OpenStudy (luigi0210):

If it's in a packet or worksheet that's fine.

OpenStudy (anonymous):

Lol...

OpenStudy (anonymous):

That's still illegitimate if he doesn't learn... People just spit answers out like it's water.

OpenStudy (luigi0210):

Besides, there could be millions of different questions out there on tests that cold be on packets and visa versa

OpenStudy (anonymous):

yuck fou

OpenStudy (anonymous):

btich

OpenStudy (anonymous):

sh is hee to ask a question ..... hey meggie write ur question in proper way thro eqation i can help u then

OpenStudy (anonymous):

\[I=\int\limits\limits_{}^{}\cot^6(x)*(\cot^2(x)+1)\csc^2(x) \rightarrow \int\limits\limits_{}^{}\cot^8(x)\csc^2(x)+\cot^6(x)\csc^2(x)\] now you can assume u=cot(x) and you will directly find that the integral equals to: \[\frac{ -\cot^9(x) }{ 9 }-\frac{ \cot^7(x) }{ 7 }\]